OpenStudy (anonymous):

Identify the correct slope and y intercept of the equation 6x - 3y = 12. slope = 2; y intercept at (0, -4) slope = -2; y intercept at (0, 4) slope = one half; y intercept at (0, 15) slope = one half; y intercept at (0, 9)

OpenStudy (anonymous):

6x-3y=12 -6x -6x -3y=12-6x divide both sides by -3 y=2x-4

OpenStudy (anonymous):

someone please help me with this

OpenStudy (anonymous):

Is it the first choice Soduko?

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

@BlveXO can you please me with that problem

OpenStudy (anonymous):

Yeah, hold on. I need help with mine first.

OpenStudy (anonymous):

okay thanks